A firm operates in a competitive market, selling good Q. His total cost is simply $200....

  1. A firm operates in a competitive market, selling good Q. His total cost is simply $200. The market price of Q is $10 a unit. If the firm is selling 96 units of Q, then what is firm’s profit?
  1. $760
  2. $960
  3. $450
  4. $120
Answer #1

Answer

Economic profit =total revenue -total cost

total revenue =P*Q=10*96=960

total cost =200

Economic profit =960-200

=$760

the profit is $760

Option 1